Using the alignment of the PZTs and BS from pre-dinner, where the beam was hitting the center of both ETMs, we aligned the DRMI. The beam was off on the SRM in yaw by ~half a beam diameter, so I undid Koji's movement of SR2 from a week ago. I loosened the SR2 dog clamps, touched it gently on the base to do a little bit of angle, then re-clamped it. Once again, Steve's new brass centering target was awesome, since it was on the SRM while I was moving SR2.
We approximately recentered the beam on the AS camera, although it didn't need much once we got the beam out of the vacuum, by centering it on all of the output AS path mirrors.
We also got IPPOS out of the vacuum. Manasa was in the process of centering the QPD when the laptop died from too long being unplugged, so we leave that for tomorrow.
Left to do:
REFL path. REFL is not coming out of the vacuum, and with the light access connector I can't reach any of the REFL steering mirrors, since they're in the center of the IOO table.
IPANG. Should be easy.
POP, POX, POY. Need to the the camera-on-a-stick back down to the corner (from ETMY) and point it at the pickoff mirrors to ensure that beam is getting out of the vacuum.